OpenStudy (anonymous):

An equation of a circle is x^2 + y^2 + 10x – 6y + 18 = 0. Show all your work in determining the center and radius of this circle. In complete sentences, explain the procedure used.

OpenStudy (abb0t):

The form of the circle is: \[(x-h)+(y-k) = r^2 \] where (h,k) is the center of the circle, and r is the radius. What you want to do is rearrange it to get it in that form. Use algebra to group (not combine) like terms. Then complete the square

OpenStudy (anonymous):

what is h and k in this equation?

OpenStudy (abb0t):

Well, that's up to you to figure out. First, group like terms: \[(x^2+10x)+(y^2-6y)+18 = 0\] Then, proceed to complete the square for x and y to get it in the form I stated earlier. Does that make more sense?

OpenStudy (abb0t):

To complete the square: \[ax^2+bx + e\] you add to both sides (including product side): \[(\frac{ b }{ 2 })^2 \]

OpenStudy (anonymous):

i know how to complete the square

OpenStudy (abb0t):

gotcha.